7.John and Janet each received an inheritance from their families. John invested his inheritance that earned an annual rate of interest of 5.25% compounded quarterly for 20 years and now had $250000 in the account. Janet invested hers at the same terms and had the same amount of money after 15 years. Algebraically calculate how much John and Janet inherited.
Answer by Boreal(15235)

m(1+(0.0525/4)^80=250000=m1.01325^80=250000=2.8388m; m=$88084.08 rounding at the end for John.
p(1+(0.0525/4)^60=200000=n(1.01325)^60=200000=2.209p; p=$90789.00 for Janet.